# Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data

### v2

@@ 5. Evaluation
We now evaluate the expressiveness of the semantic operator model as well as the efficiency of our algorithms and optimizations through four diverse applications: fact-checking (Section 5.1), extreme multi-label classification (Section 5.2), search and ranking (Section 5.3), and paper analysis (Section 5.4). We find that the semantic operator model is expressive and easy to program with, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality. We also find that our optimizations that take advantage of the declarative nature of semantic operators can accelerate semantic filtering, top-k, group-by and join operators by up to while offering statistical accuracy guarantees. Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while executing up to 28 faster end-to-end.
LOTUS’ current implementation extends the Pandas API (pan, [n. d.]). We leverage vLLM (Kwon et al., 2023) to perform efficient batched inference, and we use FAISS (Johnson et al., 2017) to support efficient vector search for LOTUS’ semantic search, similarity join and cluster operations.
@@
Unless otherwise stated, we run our local model experiments with 4 A100 GPUs using Llama-3 models (int, [n. d.]), with a batch size of 64 running on vLLM (Kwon et al., 2023). For our experiments that use OpenAI’s GPT-4o-2024-05-13 model (ope, [n. d.]), we control for cost by limiting execution to 64-way thread parallelism. We set temperature to , unless otherwise stated.
5.1. Application: Fact-Checking
@@
Application. We evaluate on FEVER (Thorne et al., 2018), a claim verification dataset. We use the development dataset, which contains about 38,000 total claims, of which we sample 1,000 for our evaluation. Each claim is labeled with one of three labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”, and the task is to correctly determine the label of each claim, leveraging evidence from a corpus of million Wikipedia articles. We merge the latter two labels in to a single class, ”Not Supported”, following prior work (Chern et al., 2023) for our evaluation.
Baselines. FacTool (Chern et al., 2023) is a recent research work that proposes a multi-step pipeline for fact-checking involving, claim extraction, query generation, tool querying, evidence collection, and verification. We use FactTool’s open source codebase (gai, [n. d.]) to measure its performance. FactTool’s pipeline, by default, performs retrieval with a Google Search API (cus, [n. d.]). We evaluate the pipeline with both the default retrieval API, and with the open-source ColBERT (Khattab and Zaharia, 2020) index. We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the implemented LOTUS programs.
LOTUS Programs. We compose several intuitive LOTUS programs, each in less than 50 lines of code. For each one, we use ColBERT as the retriever model for creating the semantic index, and we use Llama-70B as the primary LM, and Llama-8B as the cascade proxy.
First, we compose a pipeline designed to directly re-implement FacTool’s information flow in LOTUS. Figure 6 shows the pseudocode for the LOTUS-FacTool pipeline. The two tables are shown by wiki_df, which stores the Wikipedia articles, and claim_df, which stores the claims from the FEVER dataset. After loading the semantic index to wiki_df, the pipeline first performs a semantic map over each claim to generate two search queries, which are then used to perform a semantic similarity join over the corpus of Wikipedia articles. The program then concatenates the context retrieved for each claim, and performs a sem_map to output whether the claim is true or false, along with a revised claim if the claim is false. We use the same prompts found in FacTool (gai, [n. d.]), which include 3 demonstrations for generating search queries in the first sem_map, and chain-of-thought prompting in the second sem_map.
| Method | Accuracy | Execution Time (s), batched | Execution Time (s), no batching | LoC |
@@
| LOTUS-Factool | 89.9 | 688.90 | 4,454.24 | ¡ 50 |
| LOTUS-fact-filter | 91.2 | 329.1 | 988.95 | ¡ 50 |
| LOTUS-fact-filter (opt.)* | 91.0 | 189.88 | 776.37 | ¡ 50 |
| LOTUS-fact-join | 84.5 | 11,951.35 | 60,364.39 | ¡ 50 |
The next program, LOTUS-fact-filter, makes a simple, single-operator modification to the LOTUS-FacTool program, replacing the semantic map at the end of the program, with a semantic filter. We use 3 demonstrations for the filter. Figure 7 shows the pseudocode for this program. We evaluate this program with and without model cascades applied to the semantic filter.
Lastly, we compose an alternative pipeline, LOTUS-fact-join. As the pseudocode shows in Figure 8, the pipeline first performs a semantic map over each claim to obtain a set of sub-claims. From this, we create the claimed_facts_df, which separates each sub-claim into a different row. Next, the pipeline uses these sub-claims to perform a semantic similarity join over the Wikipedia corpus, then performs a semantic map over each retrieved article to generate the important facts in each one. Lastly, the program performs a semantic join between the sub-claims and the facts described in the retrieved passages. If the returned table contains a supporting fact for each sub-claim, the claim is labeled as ”Supported”. For the sem_map and join operations, we use 3 demonstrations each.
Results. Table 2 demonstrates the powerful abstraction that semantic operators provide, allowing programmers to quickly write and test programs that compose a few operators to obtain state-of-the-art results. We report the accuracy of each benchmarked method, an estimate of lines of code (LoC), and execution time in seconds both with and without batching averaged over 10 runs. We see that FacTool’s implementation offers strong accuracy performance on the FEVER dataset, however the full repository was written from scratch in several hundred lines of code, representing a significant development burden. By contrast, each LOTUS program offers comparable or higher accuracy, in relatively few lines of code. We also note that FacTool implements its method without batching, whereas LOTUS, by default, leverages batched LM execution for efficiency. To provide an apples-to-apples comparison, we compare FacTool’s un-batched implementation to the LOTUS programs both with and without batching.
First, we see from the Table 2 that LOTUS-FacTool is able to reproduce the result quality and efficiency of the original method’s implementation, with points higher accuracy and lower execution time without batching. The LOTUS-FacTool implementation with batching further decreases execution time compared to FacTool by . We see that the next LOTUS pipeline simply changes a single operation, switching the sem_map to a sem_filter, and maintains similar result quality to LOTUS-FacTool while further reducing execution time by in the batched implementation.
Leveraging LOTUS’ filter operation allows the programmer to further declaratively optimize the program using an approximation, which maintains high accuracy and reduces the batched execution time by compared to LOTUS-FacTool. Figure 9 highlights the diverse performance trade-offs presented by the model cascade optimizations used in LOTUS’ filter sem_filter. The plot shows the accuracy and execution time of the LOTUS-fact-filter pipeline using a single model, either Llama 8b, or Llama 70B,
shown by the circles. We compare this to the performance attainable using a pair of models, Llama 8B and Llama 70B, to implement the filter with our casacade-based approximation. We generate multiple cascade points, shown by the stars, by varying the recall and precision targets of the filter operator, and setting the error probability . The plot shows that this filter optimization can substantially reduce execution time, and offer diverse accuracy trade-offs, compared to implementing the pipeline with the oracle model, Llama 70B, alone. We also verify that the algorithm’s probabilistic guarantees hold by testing 100 trials for recall and precision targets at . We find the targets are met in 99 of the 100 trials, satisfying and exceeding the guaranteed 0.8 success rate.
Returning to Table 2, we find that the last LOTUS pipeline (LOTUS-fact-join), reduces accuracy and increases execution time. Notably the performance trade-offs of different AI programs are non-obvious, highlighting the need for programmable, declarative abstractions so programmers can quickly explore and iterate on their query pipelines. Each LOTUS program we described can be implemented easily in relatively few lines of code, and the best performing one offers higher accuracy and lower execution time than FacTool’s original, un-batched implementation.
5.2. Application: Extreme Multi-label Classification
@@
Dataset. We evaluate on the Biodex Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and expert-created drug safety reports constructed from each article. The task is to correctly label the drug reactions experience by the patient in each medical article. We sample 250 patient articles for our evaluation. Notably, there are possible drug-reaction labels, making this task an extreme multi-label classification task. Due to the large number of possible labels, leveraging an LM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024). We show below that this task can be efficiently modeled and programmed using the semantic join.
Baselines We consider a retrieval baseline which uses an E5Model (Wang et al., 2024) as the retriever and performs a semantic-similarity join between the patient articles and the reaction labels.
@@
| Sem-sim-join | 0.106 | 0.120 | 2.91 | 0.00 |
| LOTUS Sem-join** | 0.244 | 0.261 | 5,891.6 | 15,107 |
| nested-loop pattern | N/A | N/A | 2,144,560* | 6,092,500 |
| map-sim-filter pattern** | 0.244 | 0.261 | 5,891.6 | 15,107 |
| sim-filter pattern** | 0.154 | 0.191 | 24,206.8 | 63,724 |
@@
Estimated under linear-scaling assumption in number of batch calls.
LOTUS Programs. Our LOTUS program performs a semantic join over the drug reaction labels and the medical articles, as shown in Figure 10. We perform the semantic join using at most 7 map demonstrations. We use Llama-70b as the LM and the E5Model (Wang et al., 2024) as the retriever for the semantic index.
Results. Table 3 shows that the proposed LOTUS program makes meaningful traction, obtaining high quality results on this task, while providing query efficiency compared to naive join implementations. The table reports the rank-precision@5 (RP@5) and rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LM calls used. Since the nested-loop join pattern is prohibitively expensive to run, we show an estimate of the execution time assuming linear scaling in the number of batched calls.
First we compare performance of the LOTUS join program, declaratively optimized, to the baseline, a semantic similarity join. As expected, the optimized LOTUS join offers substantially higher quality results, with and higher RP@5 and RP@10 respectively compared to the similarity join. This highlights the effectiveness of leveraging LMs over the data for complex reasoning-based tasks. We informally compare these accuracy results to recent work (D’Oosterlinck et al., 2024) which composes a multi-step DSPy (Khattab et al., 2023) program compiled using Llama-2-7b-chat as the student LM and GPT-3.5-turbo as the teacher LM to perform a semantic mapping, followed by a re-ranking step using GPT-4-turbo. D’Oosterlinck et al. report 24.73 RP@5 and 27.67 RP@10 for the compiled program, representing comparable result quality to LOTUS’ program, although notably the LOTUS program was not compiled with a prompt optimization system.
We now consider several interesting performance trade-offs presented by semantic join algorithms. We compare the map-sim-filter join pattern, which was automatically selected in the LOTUS sem-join, to the naive nested-loop join algorithm and the sim-filter join pattern shown in Table 3. We see that nested-loop join pattern, which involves a quadratic LM budget of over 6 million LM calls, is prohibitively costly. By contrast, the sim-filter and map-sim-filter pattern substantially reduce the LM call budget of the naive algorithm by and respectively. LOTUS’ optimizer automatically selects the lower cost algorithm, in this case map-sim-filter. The selected pattern also results in significantly better accuracy due to its higher quality proxy score for this task. Specifically, the map-sim-filter pattern offers higher RP@5 and higher RP@10 compared to the sim-filter.
5.3. Application: Search & Ranking
Dataset. We evaluate LOTUS’ performance on the search and ranking task using three datasets: BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, and two new benchmarks, CIFAR-bench, and HellaSwag-bench, which we generated to evaluate more complex ranking criteria over the data. For each, we report nDCG@10 and execution time (ET) in seconds.
The SciFact dataset consists of a set of scientific claims and a corpus of articles, where the task is to rank articles by relevance to a given claim. We sample 300 claims for our evaluation, and report the average ranking execution time across these samples.
While the SciFact dataset provides a sorting task based on a simple relevance criterion, our generated benchmarks provide a more complex sorting criteria each over a corpus of paper abstracts. The task is to rank papers according to highest reported accuracy on CIFAR-10 and HellaSwag in CIFAR-bench and HellaSwag-bench, respectively.
To generate CIFAR-bench, we took 100 abstracts from the Papers with Code Dataset (pap, [n. d.]) that state performance on CIFAR-10 in the abstract, and we and manually labeled their accuracy to obtain ground truth.

### v3

@@ 5. Evaluation
We systematically evaluate LOTUS on four bulk-semantic processing applications from the AI and data literature: fact-checking, biomedical multi-label classification, search, and topic analysis. Some of these applications can be expressed by recent LLM-based analytics systems, like AI UDFs, UQE and DocETL, allowing us to compare against them; while other applications are not supported by the baseline systems, but LOTUS supports them. For each application, we additionally compare against strong hand-designed pipelines from the literature. We evaluate the LOTUS optimizer, comparing the performance of optimizations for semantic filters, joins, top-k and group-by with our gold algorithms. Lastly, we analyze the statistical accuracy guarantees provided by LOTUS’ optimizations. Overall, we find the following:
Compared to AI-based analytics systems, like AI UDFs, UQE, and DocETL, LOTUS attains similar or up to higher accuracy, while running up to faster and offering statistical accuracy guarantees that the other systems do not provide.
Compared to hand-written pipelines from the AI literature for fact-checking, biomedical classification, and search, LOTUS matches state-of-the-art quality, or exceeds it by up to , in programs involving a few semantic operators.
The LOTUS optimizer can substantially reduce cost of semantic operator programs, by up to compared to high-quality gold algorithms, and provides accuracy guarantees, which hold across repeated trials.
@@
To report controlled latency numbers, we run each baseline and experiment locally on 4 80GB A100 GPUs using Llama-3-70B (noa, [n.d.]c) and E5 embeddings (Wang et al., 2024), with a batch size of 64 running on vLLM (Kwon et al., 2023), unless otherwise stated. To run DocETL baselines and reproduce some of their benchmarks (Shankar et al., 2024), we additionally use OpenAI (noa, [n.d.]e) models, including GPT-4o-mini-2024-07-18, GPT-4o-2024-08-06 model, and text-embedding-3-small, where noted. For our experiments that use OpenAI models , we control for cost by limiting execution to 64-way thread parallelism. For reproducibility, we set temperature to for all methods and baselines, unless otherwise stated.
5.0.1. Benchmarked Methods
We briefly overview the main methods we benchmark and tested parameters.
AI UDFs. Many database vendors (ver, [n.d.]; dat, [n.d.]; noa, 2023; MotherDuck, [n.d.]; WilliamDAssafMSFT, 2024; noa, [n.d.]d) now support AI UDFs, which include map-like, row-wise LLM operations and primitives for vector search. To control for serving infrastructure, we implement these programs using LOTUS running on vLLM, and restrict our API to sem_map, sem_search and sem_sim_join.
UQE. UQE (Dai et al., 2024) proposes an optimized LLM-powered filter method, which is relevant to this work. UQE also optimizes non-LLM aggregates (e.g. COUNT) integrated with LLM-based queries, but this is beyond the scope of our evaluation. Since the code is not open-source, we implement the UQE filter in about 100 line of code in python. The UQE filter method takes an LLM call budget and provides a best-effort embedding-based approximation. In contrast, LOTUS provides a general proxy-oracle approximation method, where the proxy may be embedding-based or LLM-based. To provide a fair comparison between UQE and LOTUS, we normalize the latency budget when comparing against LOTUS programs that use LLM-based proxies, and we normalize the LLM call budget when comparing against LOTUS programs that use embedding-based proxies. We also hyper-parameter tune the mini-batch size used for UQE’s active learning algorithm.
DocETL. DocETL (Shankar et al., 2024) uses an LLM agent as the query optimizer to perform rewrites to programs, allowing programs to specify separate LLMs for the optimizer and execution. Our experiments find that the optimizer often fails, and among runs where the optimizer succeeds, performance varies. Based on personal communication with the authors, we run the DocETL baselines multiple times. We obtain three successful runs and report the average performance of these three successes. We also try using both Llama-70B as the DocETL optimizer, following our setup for each other baseline, and GPT-4o-mini as the DocETL optimizer, following the DocETL preprint (Shankar et al., 2024). However, we find both fail to produce a successful run in Section 5.2. After reaching out to the authors, we confirm a mistake in the pre-print and find that GPT-4o is required for the DocETL optimizer for our evaluation on BioDEX in Section 5.2. We also note we found several bugs in the DocETL codebase, which we have done our best to fix and have also coordinated with the authors about. Since the codebase is an evolving artifact, we report the commit number we used for our evaluation666https://github.com/ucbepic/docetl/commit/93050998077a9eb6fc1ee99dc96d1e0a222a987f.
LOTUS. We set default accuracy targets of and failure probabilities of . We provide a detailed ablation, varying these parameters in Section 5.5. Additionally, our sampling size uses of the data, similar to other works (Shankar et al., 2024), and a minimum sample size of for smaller datasets, following prior work (Kang et al., 2020).
5.1. Fact-Checking
@@
Fact-checking systems determine whether a given claim is correct, typically based on verifiability with a knowledge corpus. FacTool (Chern et al., 2023) is a recent open-source research work that provides a multi-step fact-checking pipeline, involving claim extraction, query generation, tool querying, evidence collection, and verification. We consider the task of using constructing the FacTool pipeline, which was originally written in over 750 lines of code. We use the FEVER (Thorne et al., 2018) dataset, a claim verification dataset, for our evaluation. Claim verification is based on a corpus of million Wikipedia articles, and each claim is labeled with one of three gold labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”. We merge the latter two labels into a single class, ”Not Supported”, following prior work (Chern et al., 2023) for our evaluation. To control for our API spending budget, we sample 1,000 claims from the development dataset.
Baselines. For all baselines we use ColBERT (Khattab and Zaharia, 2020) as the retreiver model777FactTool’s pipeline, by default, performs retrieval with a Google Search API (noa, [n.d.]a). We evaluate the pipeline with both the default retrieval API, and with the open-source ColBERT (Khattab and Zaharia, 2020) index. We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the other baselines. and Llama-70B served with vLLM. We run FactTool’s open source codebase (noa, [n.d.]b) to measure its performance. The AI UDF baseline uses a simple semantic map-search-map dataflow, following FacTool’s pipeline. First each claim is mapped to two search queries. Then the generated queries are used for search over the Wikipedia corpus. Lastly, each claim appended with retrieved context is mapped to truth label, along with a reasoning. We use the same prompts found in FacTool (noa, [n.d.]b), which include 3 demonstrations for generating search queries in the first semantic mapping, and chain-of-thought prompting in the second semantic mapping. The UQE baseline follows the first two steps of the AI UDF baseline, but replaces the final step with UQE’s embedding-based LLM-powered filter. Our LOTUS program similarly follows a map-search-filter pipeline and uses a Llama-8B proxy model. To provide a fair comparison, we tune the UQE LLM call budget to normalize execution time with the optimized LOTUS program. Since DocETL does not provide search primitives, we are unable to benchmark it in this application.
| Method | Accuracy | ET (s), with batching | ET (s), no batching | LoC |
@@
| AI UDF: map, search, map | 89.9 | 688.9 | 4,454.2 | ¡ 50 |
| AI UDF map, search + UQE filter | 66.0 | 184.4 | 738.3 | 150 |
| LOTUS map, search, filter (unopt.) | 91.2 | 329.1 | 989.0 | ¡ 50 |
| LOTUS map, search, filter (opt.) | 91.0 | 190.0 | 776.37 | ¡ 50 |
Results. We report the accuracy of each method, an estimate of lines of code (LoC), and average execution time over 10 runs in seconds, both with and without batching to provide a fair comparison with FacTool, which provides a sequential implementation only. Overall, Table 2 demonstrates that the optimized LOTUS program reproduces state-of-the art accuracy, comparable to that of FacTool, with faster unbatched execution and faster batched execution, and in less than 50 lines of user code, while also outperforming each baseline.
Comparing the optimized LOTUS program to the AI UDF baseline, we see that both acheive comparable accuracy in relatively few lines of code. However, the LOTUS sem_filter can be more heavily optimized compared to the generic AI UDF map operation. This allows the LOTUS program to attain faster execution, attributable to the sem_filter’s short LLM generations, which leads to cheaper decoding, and our proxy-based optimization for filters. We also observe that the LOTUS program outperforms the UQE program by accuracy at an equivalent latency budget. While our filter optimization adaptively learns when to leverage the proxy in order to meet an accuracy target, UQE’s embedding-based approximation provides no accuracy guarantees.
Finally, we compare the optimized LOTUS program to the un-optimized LOTUS program. We observe that the optimized LOTUS program achieves accuracy relative to the un-optimized one, while reducing batched execution time by . Figure 6 highlights the diverse performance trade-offs achievable by our sem_filter approximation. The plot shows the accuracy and execution time of the LOTUS program using the gold algorithm with Llama 70B, or using the proxy only, shown respectively by the green and red circles.
By varying the filter recall and precision target, our cascade-based approximation offers diverse trade-offs between accuracy and execution time, shown by the stars.
5.2. Biomedical Multi-label Classification
@@
Biomedical classification entails processing complex, lengthy patient documents. We consider the extreme multi-label classification task on the BioDEX Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and ground-truth expert-created drug safety reports constructed from each article. The task is to classify the patient’s drug reactions, given each medical article. Notably, there are possible drug-reaction labels, making this task an extreme multi-label classification task. Due to the large number of possible labels, leveraging an LLM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024). We show below that this task can be efficiently modeled and optimized using the semantic join. Similar to prior work (D’Oosterlinck et al., 2023) we sample 250 patient articles for our evaluation to control for our API spending budget.
Baselines.
The search baseline performs batched similarity search over each patient article and the set of reaction labels. For the baseline LLM-based analytics systems, we implement a LLM-powered join over the patient articles and set of candidate reaction labels, followed by a ranking step, if supported. The AI UDF program entails a naive, nested-loop join pattern using row-wise LLM operators; however we note this is prohibitively costly and we report estimated latency. While UQE does not explicitly study joins, we implement a UQE join using its optimized filter. The DocETL program uses a join and reduce, which uses listwise ranking, following the preprint (Shankar et al., 2024). We note that we are unable to successfully run DocETL using Llama-70B or GPT-4o-mini for its optimizer so results (Table 3, 4) are shown using GPT-4o for the DocETL optimizer. LOTUS supports both join and ranking. Although LOTUS provides a sem_topk operator, we limit our ranking step to list-wise ranking to maintain a fair comparison with DocETL, and revisit ranking in Section 5.3.
Our main results are shown in Table 3, using Llama-70B and E5 embeddings and sampling size of .
We also provide an additional set of benchmarks in Table 4 comparing LOTUS and DocETL following the models, prompts and parameters used in the original DocETL preprint. The DocETL setup uses GPT-4o-mini, text-embedding-3-small embeddings and a sampling size of 500. Once again, DocETL fails when we attempt to use GPT-4o-mini for the optimizer, so we use GPT-4o. DocETL’s join takes an LLM call budget, which we control to compare against the other baselines.
@@
| Search | 0.106 | 0.120 | 2.91 | 0.00 |
| AI UDF * | N/A | N/A | 2,144,560 | 6,092,500 |
| UQE | 0.115 | 0.114 | 6,559 | 15,000 |
| DocETL Join (avg of 3 successes)** | 0.180 | 0.219 | 2050 | 13,185 |
| DocETL Join + Rank (avg of 3 successes)** | 0.262 | 0.282 | 2,342 | 13,433 |
| LOTUS Join | 0.212 | 0.213 | 2,340 | 5,644 |
| LOTUS Join + Rank | 0.265 | 0.280 | 2,503 | 5,869 |
| Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
| DocETL Join + Rank (avg of 3 successes)** | 0.268 | 0.287 | 583 | 35,669 |
| LOTUS Join + Rank | 0.289 | 0.296 | 647 | 32,026 |
@@
* AI UDF is infeasible to run and latency is estimated under linear-scaling assumption in number of batched LLM calls.
** The DocETL baselines requires using GPT-4o as its optimizer and still exhibit frequent failures, with, on average, 1 in 3 runs failing. We rerun the system multiple times and report results averaged only over 3 successful runs.
Results. Table 3 demonstrates the LOTUS programs consistently match or exceed the accuracy of all baselines, while also providing performance guarantees, unlike DocETL, which requires multiple reruns and manual selection of the optimizer LLM to produce high-quality results. The table reports the rank-precision@5 (RP@5) and rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LLM calls. The LOTUS join with ranking program achieves higher rank-precision than the standalone LOTUS join program, and we show both.
We begin by informally comparing these accuracy results to that of D’Oosterlinck et al. (D’Oosterlinck et al., 2024), a state-of-the art AI pipeline that composed a multi-step DSPy (Khattab et al., 2023) program compiled using a combination of Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4-turbo. D’Oosterlinck et al. report 24.73 RP@5 and 27.67 RP@10 for the compiled program, which is comparable to result quality achieved by the simple LOTUS join with ranking program.
Turning our attention to the baselines measured in Table 3, we see that, first, the search baseline offers low results quality, with the LOTUS programs achieving over higher RP@5 and RP@10. Next, The AI UDF baseline, using row-wise LLM calls, is prohibitively expensive, and reflects the cost of our gold algorithm for joins. Its quadratic scaling of LLM call complexity with respect to dataset size leads to higher estimated cost relative to the optimized LOTUS join program. UQE, which uses an embedding-based optimization, obtains similar accuracy to the search baseline, remaining well below the RP@5 and RP@10 of the LOTUS programs despite a generous LLM call budget, indicating that the UQE method has lower sample efficiency on this task. Next, we turn to a detailed comparison of DocETL and LOTUS, shown in Table 3 and Table 4. LOTUS consistently matches or exceeds the accuracy of DocETL’s successful runs. Notably, the DocETL agentic optimizer requires GPT-4o and frequently fails, requiring multiple reruns, while LOTUS’ optimizer provides statistical accuracy guarantees, which we provide a detailed analysis on in Section 5.5. Table 4 also shows that the LOTUS and DocETL programs use a similar number of total LLM calls and execution time, however the token consumption per call of the DocETL programs are lower on average due to its optimized plan, which leads to modestly lower execution time despite modestly higher LLM calls.
| Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
| LOTUS Join Plan 1 | 0.1541 | 0.170 | 12,563 | 27,687 |
| LOTUS Join Plan 2 (chosen) | 0.212 | 0.213 | 2,116 | 5,290 |
| Gold Algorithm | N/A | N/A | 2,144,560 | 6,092,500 |
Lastly, we turn to Table 5 and study the candidate plans produced by the LOTUS join program. Plan 1, the sim-filter, requires more LLM calls to meet the recall and precision targets, compared to Plan 2, which is the project-sim-filter plan. This reflects that Plan 2’s proxy offers a stronger signal for predicate evaluations. We see that the LOTUS optimizer selects Plan 2 to execute since it is lower cost, requiring fewer LLM calls. The selected pattern also results in significantly better accuracy due to its higher quality proxy signal for this task. Specifically, the project-sim-filter pattern offers higher RP@5 and higher RP@10 compared to the sim-filter, with fewer LLM calls than the gold algorithm.
5.3. Search & Ranking